Can you help me with this double variable problem, Im not comprehending this?
4x+2y=11
-2x+5y=1
This is using the substitution method to find the x, y ..Thanks

Answer by jorel555(1290)

4x+2y=11
-2x+5y=1, or -4x+10y=2
-4x+10y=2
12y=13
y=13/12
x=53/2!!!!!!!